OpenStudy (anonymous):

the slope and line forms are eluding me any help?? write the equation of a line that contains the points (1,3) and (5,11) in slope intercept form (-2,3) (4,6) point slope form (2,0) (0,3) standard form

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) m=(11-3)/5-1) m= 8/4 m=2 y-y1=m(x-x1) y-3=2(x-1) y=2x+1 answers

OpenStudy (anonymous):

you first find the slopre between both of the points, and replace it in the slope intercept formula with x1,y1 being the ones you replace with the point closest to the origin, or if you choose the other point, change the sign for the slope